# Proactive Service Assurance in 5G and B5G Networks: A Closed-Loop Algorithm for End-to-End Network Slicing

### v3

@@ IV Proactive closed-loop algorithm design
Currently, the research on 5G network slicing is concentrated on creating a multi-service network capable of accommodating a broad spectrum of verticals, each with unique performance and service prerequisites [30]. The implementation of network slicing results in raised complexity when managing the network, particularly in scenarios involving a substantial quantity of VNF instances. Therefore, it is necessary to develop automated management and orchestration solutions. At this point, there is a lack of a cohesive consensus regarding the precise structure and extent of MANO for network slicing [5, 21, 31].
@@
When auto-scaling a VNF instance, there is a trade-off between QoS and cost saving. Within the context of network resources, the operator often allocates more resources to guarantee QoS but allocating too many resources will reduce cost savings. In network slicing, each slice is indexed by $s\in S$, where $S$ is the set of network slice instances, and it may serve multiple applications. Let $R$ be the set of compute resources, i.e., $R=\{\textsc{cpu},\textsc{ram},\textsc{sto}\}$, indexed by $r$. At any given time, each slice $s$ requires a set of VNF instances, denoted by $\text{VNFset}\textit{i}_{s}$ where $\text{VNFset}\textit{i}=\bigcup_{s\in S}{\text{VNFset}\textit{i}}_{s}$ is the set of all the VNF instances in the network. In addition, each VNF instance in VNFseti is an occurrence of a VNF, indexed by $v$. For each slice $s$, each VNF instance is allocated a specific amount of resources, i.e., cpu, ram, and sto (Storage) capacity, each denoted by $r\in R^{\textsc{$v$}}$, where $R^{\textsc{$v$}}$ defines the overall set of resources required by a VNF instance vnfi, its capacity $\textsc{cap}_{r}$ and its resource utilization (load) as ${U}_{r}$. In any part of the network from RAN to CN, there is a collection of PMs (physical machines). Note that VNF instances are instantiated in physical machines through the virtualization platform, and each of them has an amount of $\textsc{cap}_{\textsc{pm}}$. At all times, the resources utilized across all slices should not exceed those provided by the physical machines:
@@
Nowadays, 5G/B5G networks are expected to provide network connectivity to not only classical devices (i.e., tablets, smartphones, etc.) but also to the Internet of Things (IoT), which will drastically increase the number of connected devices over the network [32]. In order to assess the efficiency of a network, it is necessary to conduct regular performance monitoring. The orchestration framework often utilizes real-time monitoring, analytics, and predictive capabilities to effectively allocate network resources in accordance with traffic patterns, customer behaviour, and service demands. The implementation of dynamic resource optimization has been envisioned to improve network efficiency, minimize packet loss and reduce latency to enhance user experience. During the operation, the monitoring component will collect the network information including throughput, E2E latency, packet loss, and VNF instance information in all slices. Depending on the KPI requirements of each slice, the QoS model may vary and have different priorities. Within the context of this research, a proactive closed-loop algorithm has been proposed to operate within the CN and the TN. The algorithm is used to optimize the number of VNF instances, as well as the required cpu, ram, storage network resources, and network bandwidth, for E2E network slicing. It is assumed that:
  • All requests originating from a UE and its applications have been pre-categorized and accurately assigned to their respective slices, such as video streaming, sensitive latency applications, and so on. The implementation of this task may vary depending on the system’s design and can be accomplished through various methods [33, 34].
  • Each E2E network slice is isolated [35] by virtualization technology including network switching, router, and compute resources and can be controlled by software.
  • The scaling process is performed with a dedicated algorithm by MANO: the method of scaling may vary based on the algorithm used in MANO, e.g., creating a dedicated VNF instance with a new configuration, then switching the traffic to a new VNF instance and removing the old one. Alternatively, the existing VNF instance could be updated directly (hot scaling) during this dynamic period.
  • The monitoring component and MANO provide all the network information used in PCLANSA: network KPIs, number of VNF instances for each slice, and slice information (throughput, E2E latency, packet loss, etc.).
Figure 3: Proactive closed-loop architecture
The proposed architecture in Fig. 3 leverages the closed-loop algorithm at the slice level to enable parallel SA processing and minimize the complexity of the algorithm in development and scalability. This approach enhances the network design’s ability to expedite the processing and execution of actions. At each time window t, PCLANSA estimates the slice resource consumption per throughput unit (e.g., Mbps, Gbps) as follows:
where $\textsc{th}_{s}$ is the throughput of a given slice. By utilizing the slice resource consumption per throughput, PCLANSA is capable of efficiently calculating and identifying changes in traffic load. Therefore, it can effectively adjust resource allocation in response to fluctuations in traffic load, whether they involve an increase or decrease in resources. We can select the quantity of current VNF instances assigned to a specific network slice by taking the total resources configuration of a given slice and dividing by the maximum physical resources per VNF instance. The required number of VNF instances needed for a given slice during the upcoming time window can be computed as follows:
[…]
where $\textsc{cap}^{\max}_{sr}$ is the maximum allowed resource capacity per VNF instance that could be instantiated in a slice.
Relying upon the results derived from Formula (2), PCLANSA is able to compute the amount of resources required to facilitate the processing of a single throughput unit. Subsequently, this information can be utilized in combination with a machine learning (ML) agent model to forecast the amount of compute resources necessary for a given slice. Thus, with the assistance of a ML agent, we can predict the throughput at the next time step. This time step can be configured as $t+1$, or $t+n$, and PCLANSA can estimate the amount of compute resources and link resources $s^{r,\ell}$ needed for a given slice. Please note that $s^{r,\ell}$ is a vector formed by combining compute and link resources (creating a higher dimension vector) defined as:
$s^{r,\ell}=(\widehat{\textsc{th}}_{s}\cdot\text{Req}_{s},B^{\widehat{\textsc{th}}}_{s})$ (4)
@@
To determine the predicted throughput $\widehat{\textsc{th}}_{s}$ of a given slice, a machine learning model, LSTM-FSD, was used to analyze the historical data. Subsequently, the model was deployed in conjunction with PCLANSA to get the predicted traffic. It is important to acknowledge that the ML agent is not capable of guaranteeing 100% accuracy. Therefore, PCLANSA will implement a monitoring interval t’ to prevent abnormal traffic or incorrect prediction, and avoid excessive scaling. During the monitoring interval $t^{\prime}$, an error $e$ will be computed at each timestamp that is utilized to establish the traffic boundary. Through the implementation of this methodology, it is possible to maintain a more consistent prediction of traffic fluctuations and scaling procedures. Consequently, a traffic boundary based on a forecasting model was defined as follows:
@@
  • $\varepsilon$: is the accuracy of the traffic prediction model. For example the accuracy of the LSTM-FSD model.
@@
On the other hand, the algorithm has the capability to calculate and dynamically allocate the link resources for individual slices within the network, based on the traffic load forecast, see lines 23 to 32 in Algorithm 1 for more details. At any timestamp, the total link capacity across network slices should not exceed the link capacity provided by the network and must satisfy the equation (6).
Assuming that we have access only to the links that are connected to the core network. At any given time, each slice $s$ requires a link instance. Each link instance is allocated a specific amount of resources denoted by $\ell_{s}$, and there is also a collection of physical links $\ell^{\textsc{PHY}}\in L^{\textsc{PHY}}$. Consequently, the total link capacity allocated for network slices mustn’t exceed the total physical link capacity provided by the network infrastructure at any given time, as denoted by Equation (6):
where $\textsc{cap}_{\ell_{s}}$ is the virtual link capacity and $\textsc{cap}_{\ell^{\textsc{PHY}}}$ is the physical link capacity provided by network infrastructure. Equation (6) is utilized to verify the link configuration at any point within the network, from the RAN to the CN and from the CN to the data network, where the algorithm is executed. Finally, a proactive closed-loop algorithm, PCLANSA, was developed to address the issue of scaling in resource allocation for a specific slice. This algorithm, capable of running parallel instances across different slices, enables efficient resource allocation and scaling for network slices. By utilizing a closed-loop approach, the system can proactively allocate resources in response to changing network conditions, optimizing performance and reducing resource under-utilization. This leads to considerable enhancements in network efficiency and dependability.
@@
Input: Slice configuration parameters (Table I)
      Output: Slice configuration, network action set
1:$signal^{terminate}\leftarrow False$
2:for $\forall t$ AND $signal^{terminate}$ is $False$ do
@@
4:     $H^{Norm}_{\textsc{data}}\leftarrow pre\_process(H_{\textsc{data}})$triangleright Clean and normalized historical traffic and slice configurations.
[…]
@@
13:     if $is\_kpi\_violation$ AND $\forall Ur\leq\rho^{\textsc{ru}}$; $r\in R^{\textsc{$v$}}$  then
@@
35:     if not $enough\_resource$ then
@@
39:         Send $signal^{apply}_{resources}$ update configurations ($\text{VNFset}\textit{i}_{s}\cup s^{\ell},A$) for slice to Slice manager
@@
1:$V_{s}^{\textsc{$v$}}\leftarrow$ stack of current VNF(s) configurations for slice s
@@
3:$\rho^{\textsc{op}}_{current}\leftarrow\max\{\frac{\textsc{cap}_{s}^{current}-s^{r}}{\textsc{cap}_{s}^{current}}\}$triangleright Take the maximum over-provisioning ratio of compute resources
@@
8:$\gamma\leftarrow$ Formula (3) using $s^{r}$
9:$v\leftarrow V_{s}^{\textsc{$v$}}.pop()$
@@
11:     $R^{\textsc{comp}}\leftarrow\absolutevalue{s^{r}-\textsc{cap}_{s}^{current}}$ | $\triangleright$ $R^{\textsc{comp}}$: require resources.
@@
[…]
@@
[…]
@@
Algorithm 3 VALIDATE SLICE CONFIGURATIONS
Input: Slice configurations $s^{r,\ell}$, slice’s historical data and KPI thresholds
      Output: Validated slice configurations
1:$has\_kpi\_violation\leftarrow True$
2:while $has\_kpi\_violation$ do
3:     $has\_kpi\_violation\leftarrow$ validate $s^{r,\ell}$ by $LP_{\textsc{kpi}}$ model
@@
Definition Notation Share among slices
Accepted over-provisioning resource ratio $\rho^{\textsc{op}}$
Minimum scaling step ratio $\rho^{\textsc{s}}$
Traffic prediction accuracy ratio $\varepsilon$
Expected resource utilization ratio $\rho^{\textsc{ru}}$
Resources validation scaling ratio $\rho^{\textsc{d}}$
Maximum allocated resources per VNF instance for slice $\textsc{cap}^{\max}_{sr}$
Minimum allocated resources per VNF instance for slice $\textsc{cap}^{\min}_{sr}$
Number of sampling data $\kappa$
Physical nodes configuration $\textsc{cap}_{\text{PMset}}$
Total physical link capacity $L$
Time window $t$
Monitoring traffic time window $t^{\prime}$
Set of target slice KPIs $KPI^{s}$
@@
Core Network 1 $scale\_up$ Scale-up slice resources.
2 $scale\_down$ Scale-down slice resource.
3 $scale\_out$ Add VNF instance(s).
4 $scale\_in$ Remove VNF instance(s).
Transport Network 5 $scale\_up\_link$ Increase virtual link capacity.
6 $scale\_down\_link$ Decrease virtual link capacity.
Both 7 $no\_action$ No action need.
@@
Figure 4: Overview of PCLANSA: a high-level flowchart.
@@
The high-level flowchart in Fig. 4 illustrates how our closed-loop algorithm works. A closed-loop algorithm in network slicing aims to provide an efficient and dynamic service deployment and management capable of maintaining the QoS of multiple slices by detecting KPI violations quickly and accurately, taking corrective actions, and assigning appropriate resources to resolve KPI violations in a timely manner. Thus, the algorithm has been developed with flexible parameters, shown in Table I, enabling operators to customize PCLANSA by themselves in order to meet the specific requirements of the slice and align it to their infrastructure. To keep things simple, we split PCLANSA into two main algorithms:
@@
In the first Algorithm 1, the algorithm aims to forecast the upcoming traffic (per slice), see lines 3 to 6. By combining historical information with equations (2) and (4), the algorithm estimates the resources required for the forthcoming timestamp, as shown in lines 7 to 11. Before optimizing resources and performing scaling, the algorithm conducts a KPI violation check in line 12. If all resources are below the $\rho^{\textsc{ru}}$ rate but have KPI violations, there is a possibility of an abnormal event in the system, such as a dropped link connection or loss of power in a node. In such cases, the algorithm will trigger an alarm in the system. In the event of a KPI violation, the algorithm will attempt to increase resources by a factor of $\rho^{\textsc{op}}$ to mitigate potential bottlenecks caused by insufficient resources. Subsequently, the configuration of the slice is evaluated using the $LP_{\textsc{kpi}}$ model to estimate the necessary resources, with the objective of preventing KPI violations associated with the slice. Afterwards, the final slice configuration will be processed by Algorithm 3 which will perform precise network checks. The algorithm verifies the network infrastructure constraints and ensures that there is sufficient network capacity for the slice before executing any operations.
@@
The second Algorithm 2 is used to optimize the slice configuration obtained in the first phase and perform accurate actions, shown in Table II. To mitigate the issue of frequent scaling, the algorithm utilizes a minimum scaling increment denoted by $\rho^{\textsc{s}}$. It is used to determine the minimum quantity of resources required in the event of infrastructure expansion. During the process of scaling up or scaling down, the algorithm allocates resources for each compute resource type independently, in order to ensure that each resource type is in accordance with the upcoming traffic. In the scaling out phase, the algorithm tries to maximize the resources of the last VNF while maintaining a consistent ratio of values among compute resources. Subsequently, the algorithm computes and adds a new VNF into the slice. The aforementioned mechanism is also applicable to the scaling in phase but in the reverse direction. The algorithm computes the optimal resources required for VNFs in the subsequent period while aligning it with the slice KPIs. More details on LSTM-FSD and $LP_{\textsc{kpi}}$ models can be found in our previous work [36].
While the closed-loop algorithm for E2E network slicing in 5G/B5G networks is promising, there are several potential limitations and challenges associated with implementing it in a real-world network. Some of these challenges may include:
  • Data availability and quality: Accurate prediction of the algorithm is based on the availability and quality of data relating to network traffic and resource utilization. In the context of a real 5G network, the task of gathering and analyzing data in real-time may present difficulties, potentially impacting the precision of the prediction phase in the algorithm, especially E2E data.
  • Computational complexity: The algorithm integrates machine learning and linear programming methodologies, which can result in significant computational demands. The computational complexity of the algorithm in a network with numerous VNFs and slices may pose a significant challenge, potentially impacting its performance and scalability.
  • Parameter tuning: The algorithm has a total of 13 configurable parameters, which need to be carefully tuned in order to optimize its performance. In a real-world network, finding the optimal values for these parameters could be challenging, as it may require extensive experimentation and testing.
  • Adaptability to changing conditions: 5G/B5G networks are dynamic and can experience rapid changes in traffic patterns and resource requirements. Therefore, the algorithm needs to adapt quickly to these changing conditions to continue making accurate predictions and allocating resources effectively. Consequently, enabling the algorithm to promptly adapt to real-world network fluctuations may pose a challenge.
The next part of this paper will present a comprehensive experiment utilizing PCLANSA and demonstrate our research results.

### v4

@@ IV Proactive closed-loop algorithm design
IV-A Resource model
@@
Each network slice $s$ ( $s\in S$, where $S$ is the set of network slice instances) can utilize multiple VNF instances $v$ ($\textsc{$v$}\in\text{VNFset}\textit{i}_{s}$, where $\text{VNFset}\textit{i}_{s}$ is the set of VNF instances for slice $s$). Each VNF instance $v$ requires resources $r$ ($r\in R=\{\textsc{cpu},\textsc{ram},\textsc{sto}\}$ with capacity $\textsc{cap}_{r}^{\textsc{$v$}}$ and untilization $U_{r}^{\textsc{$v$}}$. In the network, there is a set of physical machines ($\textsc{PM}s$) that have resource capacities $\textsc{cap}_{r}^{\textsc{pm}}$, where $\textsc{pm}\in\textsc{PM}s$. Note that VNF instances are instantiated in physical machines through the virtualization platform, and each of them has an amount of $\textsc{cap}^{\textsc{pm}}$. At all times, the resources utilized across all network slices should not exceed those provided by the physical machines:
$\sum_{s\in S}\sum_{\textsc{$v$}\in\text{VNFset}\textit{i}_{s}}\textsc{cap}^{\textsc{$v$}}_{r}\leq\sum_{\textsc{pm}\in\text{PMset}}\textsc{cap}^{\textsc{pm}}_{r},\forall r\in R$ (1)
At any timestamp, the total link capacity across network slices should not exceed the link capacity provided by the network and must satisfy the equation (2). Assuming that we have access only to the links that are connected to the core network. At any given time, each network slice $s$ requires a link instance. Each link instance is allocated a specific amount of resources denoted by $\ell_{s}$, and there is also a collection of physical links $\ell^{\textsc{PHY}}\in L^{\textsc{PHY}}$. Consequently, the total link capacity allocated for network slices mustn’t exceed the total physical link capacity provided by the network infrastructure at any given time, as denoted by Formula (2):
where $\textsc{cap}_{\ell_{s}}$ is the virtual link capacity and $\textsc{cap}_{\ell^{\textsc{PHY}}}$ is the physical link capacity provided by network infrastructure. Formula (2) is utilized to verify the link configuration at any point within the network, from the RAN to the CN and from the CN to the data network, where the algorithm is executed.
IV-B Implementation of PCLANSA
Figure 3: Proactive closed-loop architecture
The proposed architecture in Fig. 3 leverages the closed-loop algorithm at the network slice level to enable parallel SA processing and minimize the complexity of the algorithm in development and scalability. This approach enhances the network design’s ability to expedite the processing and execution of actions. At each time window t, PCLANSA estimates the network slice resource consumption per throughput unit (e.g., Mbps, Gbps) as follows:
where $\textsc{th}_{s}$ is the throughput of a given network slice. By utilizing the network slice resource consumption per throughput, PCLANSA is capable of efficiently calculating and identifying changes in traffic load. Therefore, it can effectively adjust resource allocation in response to fluctuations in traffic load, whether they involve an increase or a decrease in resources. We can select the quantity of current VNF instances assigned to a specific network slice by taking the total resources configuration of a given network slice and dividing by the maximum physical resources per VNF instance. The required number of VNF instances needed for a given network slice during the upcoming time window can be computed as follows:
$\gamma_{s}=\left\lceil\max\limits_{R}\left\{\frac{\sum\limits_{\textsc{$v$}\in\text{VNFset}\textit{i}_{s}}\sum\limits_{\textsc{vnf}\textit{i}}\sum\limits_{r\in R^{\textsc{$v$}}}\textsc{cap}_{r}^{\textsc{vnf}\textit{i}}}{\textsc{cap}^{\max}_{sr}}\right\}\right\rceil$ (4)
where $\textsc{cap}^{\max}_{sr}$ is the maximum allowed resource capacity per VNF instance that could be instantiated in a network slice.
Relying upon the results derived from Formula (3), PCLANSA is able to compute the amount of resources required to facilitate the processing of a single throughput unit. Subsequently, this information can be utilized in combination with a machine learning (ML) agent model to forecast the amount of compute resources necessary for a given network slice. Thus, with the assistance of an ML agent, we can predict the throughput at the next time step. This time step can be configured as $t+1$, or $t+n$, and PCLANSA can estimate the amount of compute resources and link resources $s^{r,\ell}$ needed for a given network slice. Please note that $s^{r,\ell}$ is a vector formed by combining compute and link resources (creating a higher dimension vector) defined as:
@@ IV-B Implementation of PCLANSA
To determine the predicted throughput $\widehat{\textsc{th}}_{s}$ of a given network slice, a machine learning framework, $LP_{\textsc{kpi}}$ [29], was used to analyze the historical data. The $LP_{\textsc{kpi}}$ framework comprises two components: the LSTM-FSD model, which performs short-term throughput forecasting using traffic, resource utilization, and network slice configuration data; and the LP-KPI model, which employs an ILP-based approach to predict additional KPIs, such as delay and packet loss, by integrating the predicted throughput with the current network state. Subsequently, the model was deployed in conjunction with PCLANSA to get the predicted traffic and estimate network KPIs in the upcoming time. It is important to acknowledge that the ML agent is not capable of guaranteeing 100% accuracy. Therefore, PCLANSA will implement a monitoring interval $t^{\prime}$ to prevent abnormal traffic or incorrect prediction, and avoid excessive scaling. During the monitoring interval $t^{\prime}$, an error rate $E_{t^{\prime}}$ will be computed at each timestamp that is utilized to establish the traffic boundary. This error rate can be calculated by determining the mean of the absolute differences between the predicted and actual throughput for each time step within the monitoring window, $t^{\prime}$. Through the implementation of this methodology, it is possible to maintain a more consistent prediction of traffic fluctuations and scaling procedures. Consequently, a traffic boundary based on a forecasting model was defined as follows:
@@
  • $\varepsilon$: is the accuracy of the traffic prediction model. For example, the accuracy of the LSTM-FSD model.
@@
In addition, the algorithm possesses the capacity to calculate and dynamically allocate link resources for individual network slices within the network, contingent on the traffic load forecast (refer to lines 23 to 32 in Algorithm 1 for further details). The specifics of our PCLANSA are delineated in Algorithm 1. This algorithm is designed to run parallel instances across different network slices, thereby enabling efficient resource allocation and scaling for network slices. Utilizing a closed-loop with the ML approach, the system can proactively allocate resources in response to changing network conditions, thereby optimizing performance and reducing resource under-utilization. Thus, this leads to considerable enhancements in network efficiency and dependability.
Figure 4: Overview of PCLANSA: a high-level flowchart.
@@
Input: Network slice configuration parameters (Table I)
   Output: Network slice configuration, network action set
[…]
2:for $\forall t$ AND $signal^{terminate}$ is FalseFalse do
@@
4:  $H^{Norm}_{\textsc{data}}\leftarrow pre\_process(H_{\textsc{data}})$ $\triangleright$ Clean and normalized historical traffic and network slice configurations.
5:  $\widehat{\textsc{th}}\leftarrow\textit{LSTM-FSD}$ agent by $H^{Norm}_{\textsc{data}}$
@@
13:  if is_kpi_violationis\_kpi\_violation AND $\forall Ur\leq\rho^{\textsc{ru}}$; $r\in R^{\textsc{$v$}}$ then
@@
35:  if not enough_resourceenough\_resource then
@@
39:   Send $signal^{apply}_{resources}$ update configurations ($\text{VNFset}\textit{i}_{s}\cup s^{\ell},A$) for network slice to Network Slice manager
@@
[…]
@@
[…]
@@
11:  $|R^{\textsc{comp}}\leftarrow\absolutevalue{s^{r}-\textsc{cap}_{s}^{current}}$ $\triangleright$ $R^{\textsc{comp}}$: require resources.
@@
26:  $R^{new}\leftarrow\max\{\textsc{cap}^{\min}_{sr},R^{\textsc{comp}}-\sum\limits_{\textsc{$v$}\in V_{s}^{\textsc{$v$}}}\sum\limits_{r\in R^{\textsc{$v$}}}\sum\limits_{\textsc{vnf}\textit{i}}\textsc{cap}_{r}^{\textsc{vnf}\textit{i}}\}$ $\triangleright$ $R^{new}$: resources for new VNF instance $v^{new}$.
@@
29:  $p\leftarrow min\{\frac{\sum\limits_{\textsc{$v$}\in V_{s}^{\textsc{$v$}}}\sum\limits_{r\in R^{\textsc{$v$}}}\sum\limits_{\textsc{vnf}\textit{i}}\textsc{cap}_{r}^{\textsc{vnf}\textit{i}}-s^{r}}{\sum\limits_{r\in R^{V_{s}^{\textsc{$v$}}[Last]}}\textsc{cap}^{\textsc{vnf}\textit{i}}_{r}}\}$
@@
Algorithm 3 VALIDATE NETWORK SLICE CONFIGURATIONS
Input: Network slice configurations $s^{r,\ell}$, slice’s historical data and KPI thresholds
   Output: Validated network slice configurations
[…]
2:while has_kpi_violationhas\_kpi\_violation do
3:  $has\_kpi\_violation\leftarrow$ validate $s^{r,\ell}$ by $LP_{\textsc{kpi}}$ framework
@@
Notation Definition
Share
among network slices
$\rho^{\textsc{op}}$ Accepted over-provisioning resource ratio
$\rho^{\textsc{s}}$ Minimum scaling step ratio
$\varepsilon$ Traffic prediction accuracy ratio
$\rho^{\textsc{ru}}$ Expected resource utilization ratio
$\rho^{\textsc{d}}$ Resources validation scaling ratio
$\textsc{cap}^{\max}_{sr}$ Maximum allocated resources per VNF instance for network slice
$\textsc{cap}^{\min}_{sr}$ Minimum allocated resources per VNF instance for network slice
$\kappa$ Number of sampling data
$\textsc{cap}_{\text{PMset}}$ Physical nodes configuration
$L$ Total physical link capacity
$t$ Time window
$t^{\prime}$ Monitoring traffic time window
$KPI^{s}$ Set of target network slice KPIs
@@
Core Network 1 scale_upscale\_up Scale-up network slice resources.
2 scale_downscale\_down Scale-down network slice resource.
3 scale_outscale\_out Add VNF instance(s).
4 scale_inscale\_in Remove VNF instance(s).
Transport Network 5 scale_up_linkscale\_up\_link Increase virtual link capacity.
6 scale_down_linkscale\_down\_link Decrease virtual link capacity.
Both 7 no_actionno\_action No action needed.
@@
As illustrated in the high-level flowchart in Figure 4, the PCLANSA functions by leveraging the aforementioned formulas embedded within the ML model. Our proactive closed-loop algorithm in the network aims to provide an efficient and dynamic service deployment and management capable of maintaining the QoS of multiple network slices by detecting KPI violations quickly and accurately, taking corrective actions, and assigning appropriate resources to resolve KPI violations on time. Thus, the algorithm has been developed with flexible parameters, shown in Table I, enabling operators to customize PCLANSA by themselves in order to meet the specific requirements of the network slice and align it to their infrastructure. To keep things simple, we split PCLANSA into two main algorithms:
@@
In the first Algorithm 1, the algorithm aims to forecast the upcoming traffic (per network slice), see lines 3 to 6. By combining historical information with Formulas (3) and (5), the algorithm estimates the resources required for the forthcoming timestamp, as shown in lines 7 to 11. Before optimizing resources and performing scaling, the algorithm conducts a KPI violation check in line 12. If all resources are below the $\rho^{\textsc{ru}}$ rate but have KPI violations, there is a possibility of an abnormal event in the system, such as a dropped link connection or loss of power in a node. In such cases, the algorithm will trigger an alarm in the system. In the event of a KPI violation, the algorithm will attempt to increase resources by a factor of $\rho^{\textsc{op}}$ to mitigate potential bottlenecks caused by insufficient resources. Subsequently, the configuration of the network slice is evaluated using the $LP_{\textsc{kpi}}$ framework to estimate the necessary resources, with the objective of preventing KPI violations associated with the network slice. Afterwards, the final network slice configuration will be processed by Algorithm 3, which will perform precise network checks. The algorithm verifies the network infrastructure constraints and ensures that there is sufficient network capacity for the network slice before executing any operations.
@@
The second Algorithm 2 is used to optimize the network slice configuration obtained in the first phase and perform accurate actions, shown in Table II. To mitigate the issue of frequent scaling, the algorithm utilizes a minimum scaling increment denoted by $\rho^{\textsc{s}}$. It is used to determine the minimum quantity of resources required in the event of infrastructure expansion. During the process of scaling up or scaling down, the algorithm allocates resources for each compute resource type independently, in order to ensure that each resource type is in accordance with the upcoming traffic. In the scaling-out phase, the algorithm tries to maximize the resources of the last VNF (in the same network slice) while maintaining a consistent ratio of values among compute resources. Subsequently, the algorithm computes and adds a new VNF into the network slice only if the last VNF has exhausted its maximum allowable resources. The aforementioned mechanism is also applicable to the scaling in phase but in the reverse direction. Consequently, the algorithm is capable of calculating the optimal resources necessary for VNFs in the subsequent period, aligning them with the network slice KPIs.